OpenStudy (dinorap1):

Please help! Will give medal! "A fox sees a piece of carrion being thrown from a hawk's nest and rushes to snatch it. The nest is 14.0 m high, and the carrion is thrown with a horizontal velociity of 1.5 m/s. The fox is 7.0 m from the base of the tree. What is the magnitude of the fox's average velocity if it grabs the carrion in its mouth just as it touches the ground? (Assume no air resistance and that a(y)=g=9.81m/s^2) c. 4.2 m/s d. 2.6 m/s" (My teacher told me that the answer to this question was either c or d)

OpenStudy (anonymous):

Here is what I have, the nest is 14m above the ground, we can solve for time knowing this, using yf=1/2at^2+v(initial)t+y(i), we input -14=1/2(-9.8(for gravity)t^2+0(t)+0 equals -14=-4.9t^2,,, -14/-4.9=t^2. t=+/-1.69

OpenStudy (anonymous):

now that we know t we can input it for the x values. xf=1/2(a)t^2+v(i)t+x(i). xf=-4.99(0)1.69)^2+1.5(1.69)+0 xf=2.53

OpenStudy (anonymous):

since we know the fox was 7 meters away, we subtract the distance for the x value which was 2.53 from 7. 7-2.53=4.47m. this is the distance the fox has to travel in the 1.69 seconds it takes the carrion to be 2.53 m away from the nest

OpenStudy (anonymous):

we take the 4.47/1.69=2.64m/s

OpenStudy (dinorap1):

What are the formulas you're using to figure this out? Sorry, I wish I was here when you were writing all this...

OpenStudy (anonymous):

What we know for the carrion falling: X=-14m, a=-9.8m/s^2, V_0=0m/s. So I will use this equation\[X=V_0t+1/2at^2\] \[-14m=(0m/s)(t)+1/2(-9.8m/s^2)(t^2)\] t=1.69s I assume the carrion is being thrown at the fox? Anyways 1.5m/s*1.69s = 2.535m so that's how far the carrion moves toward the fox so 7m-2.535m = 4.465m is the total distance the fox must move in the 1.69s. \[4.465m/1.69s=2.64m/s\] I hope I broke it down enough! Any questions please ask.